OpenStudy (anonymous):

how do i evaluate dy when dy = sec^2(pi/3)(-.05) ?

OpenStudy (turingtest):

I believe you are missing some info here...

OpenStudy (anonymous):

what do you mean

OpenStudy (turingtest):

y=sec^2(pi/3)(-.05x) perhaps?

OpenStudy (turingtest):

not dy=... then you would have nothing to do

OpenStudy (mathteacher1729):

Do you know how to evaluate \(\sin(\pi/3)\) and \(\cos(\pi/3)\) ? Remember that \(\sec(x) = \frac{1}{\cos(x)}\).

OpenStudy (anonymous):

no i already plugged in values for x and dx where dx is -.05 and x is pi/3

OpenStudy (mathteacher1729):

The question is asking to *evaluate* -- so there is nothing that needs to be solved. If you know what cos(pi/3) is... take 1 over cos(pi/3) , square that, and multiply it by -1/2. :)

OpenStudy (anonymous):

hmm..im confused haha

OpenStudy (anonymous):

dy = Sec[Pi/3]^2 - 0.05 = 3.95